a student is only able to read the lecture notebook when it is held at 65.00cm from his eyes. what focal length and power must the lenses of his reading glasses have to bring his near point to the normal value


1
Expert's answer
2022-02-10T14:53:58-0500

"F=\frac{dd_0}{d-d_0}=40.6~cm,"

"D=\frac 1F=2.46~dpt."